Circular No. 8077

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003av AND 2003aw Further to IAUC 8071, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and K. Li report the discovery of two apparent supernova on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003av Feb. 4.28 8^{h}01^{m}32.^{s}57 + 2^{o}48′27′′.9 19.6 3′′.5 E, 1′′.2 N
2003aw Feb. 6.26 9 05 54.79 − 5 36 08.6 17.8 0′′.2 E, 1′′ N

Additional magnitudes: SN 2003av, Jan. 16 UT, [21.0; Feb. 10, 19.5; 18, 19.7; 19, 20.0. SN 2003aw, Jan. 14, [21.0; Feb. 10, 17.6; 19, 17.8.

IGR J16320–4751 AND IGR J16318–4848 J. J. M. in ’t Zand, Space Research Organization Netherlands; and P.

Ubertini, F. Capitanio, and M. Del Santo, Istituto Astrofisica Spaziale e Fisica Cosmica, Rome, report on their analysis of archival BeppoSAX-WFC data. IGR J16320–4751 = AX J1631.9–4752 (IAUC 8076) was consistently detected during 1996–2002, at a flux that varied between 10 and 16 mCrab (5–10 keV; 2-month averages); with the ASCA detection, this suggests that the source has been persistently active for at least 8 years. Based on three broad channels (2–5, 5–10, and 10–28 keV) and averaged over all data, an absorbed power model describes the data well with a photon index of 2.5 ± 0.3 and N_{H} = (2 ± 1) × 10^{23} cm^{−2}. The average unabsorbed 0.7–10-keV flux is 4 × 10^{−10} erg cm^{−2} s^{−1}. If the source is a high-mass x-ray binary (cf. IAUC 8076), the persistently bright emission would be more in line with the mass donor being a (super)giant than a Be star. The position in the combined WFC data is α = 16^{h}32^{m}05^{s}.4, δ = −47^{o}52′07′′ (equinox 2000.0; 99-percent-confidence error radius 1′.7), which is 1′.1 from the centroid of AX J1631.9–4752 and 1′ from IGR J16320–4751. IGR J16318–4848 (IAUC 8063, 8076) was never detected in any accumulation of WFC data, with an upper limit of 5 mCrab per visibility window.

R CORONAE BOREALIS J. Mattei, AAVSO, reports that R CrB last faded in late Nov. 2000,

reaching minimum m_{vis} = 12.4 (mean maximum m_{vis} = 6.3 in Feb. 2001). Visual magnitude estimates: 2003 Feb. 4.14 UT, 6.2 (A. Baransky, Bucha, Ukraine); 11.12, 6.6 (Baransky); 13.98, 7.0 (K. Hornoch, Lelekovice, Czech Rep.); 15.02, 7.2 (Hornoch); 17.11, 7.8 (Baransky); 21.12, 9.1 (Baransky).

2003 February 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green